### v1

# Quench dynamics via recursion method and dynamical quantum phase transitions

Abstract
The recursion method provides a powerful framework for studying quantum many-body dynamics in the Lanczos basis, which is recursively constructed within the Krylov space of operators. Recently, it has been demonstrated that the recursion method, when supplemented by the universal operator growth hypothesis, can effectively compute autocorrelation functions and transport coefficients at infinite temperature. In this work, we extend the scope of the recursion method to far-from-equilibrium quench dynamics. We apply the method to spin systems in one, two, and three spatial dimensions. In one dimension we find a clear evidence that the convergence of the method is limited in time by the dynamical quantum phase transition.
Introduction.  Addressing quantum dynamics by means of the recursion method has a long history Mori (1965); Dupuis (1967); Haydock (1980); Mattis (1981); Joslin and Gray (1986); Roldan et al. (1986); Viswanath and Müller (2008); Lindner and Auerbach (2010); Nandy et al. (2024). In simple words, the (Heisenberg version of the) method amounts to constructing an orthonormal Lanczos basis within the Krylov space of an evolving quantum observable and harnessing the simplifications this basis brings about. A new impact has been given to the method by the universal operator growth hypothesis Parker et al. (2019), which has informed about the generic pattern of the evolution operator in the Lanczos basis.
Up to now the practical application of the recursion method to quantum many-body systems has been limited almost exclusively by computing infinite-temperature autocorrelation functions and transport coefficients Parker et al. (2019); Khait et al. (2016); de Souza et al. (2020); Yates et al. (2020a, b); Yuan et al. (2021); Wang et al. (2024); Bartsch et al. (2024); Uskov and Lychkovskiy (2024); Teretenkov et al. (2024); Bhattacharyya et al. (2024). There the method has demonstrated high efficiency. This efficiency is particularly notable in application to two-dimensional systems Uskov and Lychkovskiy (2024); Teretenkov et al. (2024); Bhattacharyya et al. (2024), where traditional numerical methods often struggle. Another notable advantage of the recursion method is that it allows a straightforward symbolic implementation, when a single computation covers all parameters of the Hamiltonian Uskov and Lychkovskiy (2024); Teretenkov et al. (2024).
Formally, the output of the recursion method is the Heisenberg operator of an observable. This suggests that the scope of the method can be much broader than infinite-temperature correlation functions and should cover, in particular, the out-of-equilibrium dynamics after quantum quenches. Addressing the later in systems of dimension higher than one is known to be highly challenging, with limited body of results obtained to date Schmitt and Heyl (2020); Dziarmaga (2022); Bernier and Agarwal (2025); Begušić and Chan (2024); Loizeau et al. (2025a). Given the success of the recursion method in computing two-dimensional autocorrelation function, it seems tempting to try to apply it to quench dynamics.
Here we undertake exactly this attempt. The paper is organized as follows. We start from briefly introducing the recursion method and specifying it for addressing quench dynamics. Then we apply it to a paradigmatic one-dimensional mixed-field Ising model. The results are benchmarked by essentially exact numerical simulations. We demonstrate that the recursion method has a convergence domain limited in time by the dynamical quantum phase transition. Then some results for the transverse-field Ising model on the square and cubic lattices are presented. We conclude by discussing implications of our results and open questions posed by them.
Recursion method for quench dynamics.  Consider an observable given by a self-adjoint Shrödinger operator $A$. The same observable in the Heisenberg representation reads $A_{t}=e^{itH}A\,e^{-itH}$, where $H$ is the Hamiltonian of the system. The Heisenberg operator $A_{t}$ evolves within the Krylov space spanned by operators $A,{\mathcal{L}}A,{\mathcal{L}}^{2}A,\dots$, where ${\mathcal{L}}\equiv[H,\bullet]$ is the superoperator referred to as Liouvillian. $A_{t}$ satisfies the Heisenberg equation of motion $\partial_{t}A_{t}=i\,{\mathcal{L}}A_{t}$. Within the recursion method Viswanath and Müller (2008), this equation is addressed in the orthonormal Lanczos basis $\{O^{0},O^{1},O^{2},\dots\}$ constructed as follows: $O^{0}=A/\|A\|$, $A^{1}={\mathcal{L}}O^{0}$, $b_{1}=\|A^{1}\|$, $O^{1}=b_{1}^{-1}\,A^{1}$ and $A^{n}={\cal L}\,O^{n-1}-\,b_{n-1}O^{n-2}$, $b_{n}=\|A^{n}\|$, $O^{n}=b_{n}^{-1}A^{n}$ for $n\geq 2$. Here a scalar product in the space of operators is introduced according to $\big{(}A|B\big{)}\equiv\operatorname{\mbox{tr}}\big{(}A^{\dagger}B\big{)}/d$, where $d$ is the Hilbert space dimension (which is assumed to be finite). The scalar product entails the norm $\|A\|=\sqrt{(A|A)}$. The coefficients $b_{n}$ are referred to as Lanczos coefficients. In this basis the Liouvillian has the tridiagonal form,
${\mathcal{L}}=\begin{pmatrix}0&b_{1}&0&0&\dots\\[3.00003pt] b_{1}&0&b_{2}&0&\dots\\[3.00003pt] 0&b_{2}&0&b_{3}&\dots\\[3.00003pt] 0&0&b_{3}&0&\dots\\[3.00003pt] \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}.$ (1)
The solution of the Heisenberg equation for the observable $A_{t}$ can be expressed as
$A_{t}=\sum_{n=0}^{\infty}\phi^{n}_{t}\,O^{n},$ (2)
where $\phi^{n}_{t}\equiv(O^{n}|A_{t})$ are the solution of the system of coupled equations, which can be written in the matrix form as
$\partial_{t}\phi_{t}=$ $i\,{\mathcal{L}}\phi_{t}.$ (3)
Here $\phi_{t}=\|\phi^{n}_{t}\|_{n=0,1,2,\dots}$ is a column vector and ${\mathcal{L}}$ is the matrix (1).
In practice, one can explicitly compute only a finite number $n_{\rm max}$ of Lanczos coefficients. The rest of the coefficients should be extrapolated. The universal operator growth hypothesis Parker et al. (2019) (see also precursor work Liu and Müller (1990); Florencio et al. (1992); Zobov and Lundin (2006); Elsayed et al. (2014); Bouch (2015)) states that in a generic quantum many-body system Lanczos coefficients $b_{n}$ scale asymptotically linearly with $n$ at large $n$ (with a logarithmic correction for one-dimensional systems). This asymptotic behavior has been confirmed in various many-body models Cao (2021); Noh (2021); Heveling et al. (2022); Uskov and Lychkovskiy (2024); De et al. (2024). As a rule, the leading asymptotics can be easily extracted from the first dozen of Lanczos coefficients. However, extrapolating Lanczos coefficients by a linear function is a too rough approximation: the qualitative behavior of the solutions of the Heisenberg equation strongly depends on the subleading asymptotic terms Viswanath et al. (1994); Parker et al. (2019); Yates et al. (2020a, b); Dymarsky and Smolkin (2021); Bhattacharjee et al. (2022); Avdoshkin et al. (2024); Camargo et al. (2023); Uskov and Lychkovskiy (2024); Dodelson (2025). Guided by the results of ref. Gam , we employ the following extrapolation formula:
$b_{n}$ $\simeq\frac{\alpha\,n}{\log n}+\gamma+\frac{(-1)^{n}\alpha}{\log n+c}$ ${\rm for}~{}~{}1{\rm D},$ (4)
$b_{n}$ $\simeq\alpha\,n+\gamma+\frac{(-1)^{n}\alpha}{\log n+c}$ ${\rm for}~{}~{}2,3{\rm D}.$ (5)
The last alternating term is required whenever $\varphi^{0}_{t}$ relaxes to a nonzero value at $t\rightarrow\infty$ Gam , which is generally the case. Parameters $\alpha,\,\gamma$ and $c$ are found by fitting $n_{\rm max}$ explicitly computed Lanczos coefficients.
The knowledge of the Heisenberg operator $A_{t}$ of the observable allows one to address the quench dynamics, i.e. compute the time-dependent value $\langle A_{t}\rangle\equiv\operatorname{\mbox{tr}}\rho_{0}A_{t}$ of the observable that evolves after the system is initialized in an out-of-equilibrium state $\rho_{0}$. Within our approach, we compute $\langle A_{t}\rangle$ according to
$\langle A_{t}\rangle\simeq\sum_{n=0}^{n_{\rm max}}\phi^{n}_{t}\,\langle O^{n}\rangle,$ (6)
where $\langle O^{n}\rangle\equiv\operatorname{\mbox{tr}}\rho_{0}O^{n}$. We compute $O^{n}$ and $b_{n}$ for $n$ up to $n_{\rm max}$ symbolically directly in the thermodynamic limit, as described in Uskov and Lychkovskiy (2024). $\langle O^{n}\rangle$ is also computed symbolically. As for functions $\phi^{n}_{t}$, they are obtained by numerically solving coupled linear differential equations (3) with the tridiagonal matrix ${\mathcal{L}}$ truncated at some large $k$ (we use $k=1000$ in the examples below). The entries of this matrix for $n>n_{\rm max}$ are obtained from extrapolation of the $n_{\rm max}$ known Lanczos coefficients according to extrapolation formulae (4), (5).
Note that $\|A\|\,\phi^{0}_{t}=d^{-1}\,\operatorname{\mbox{tr}}A\,A_{t}$ is the infinite-temperature autocorrelation function of the observable $A$. Recently the recursion method supplemented by the UOGH has proven effective in computing such autocorrelation functions Parker et al. (2019); Wang et al. (2024); Uskov and Lychkovskiy (2024); Teretenkov et al. (2024). The out-of-equilibrium post-quench expectation value $\langle A_{t}\rangle$ is, however, a more complex object. Below we explore it in specific models.
Figure 1: Time evolution of the (a) Loschmidt rate and (b)–(d) total polarization along $z$ direction in the 1D Ising model (7) with $h_{x}=h_{z}=1$. The initial state is a product translation-invariant state (9) with the polarization $(p_{x},p_{y},p_{z})$ indicated on each plot. Orange line – essentially exact results of the numerical time evolution (see text for details). Blue line - results of the recursion method. The Loschmidt rate features multiple cusp singularities manifesting dynamical quantum phase transitions. The results of the recursion method depart from the exact numerical results at time $t_{*}$ close to the time $t_{c}$ of the second cusp (red dashed vertical line).
Figure 2: Scaling of the estimated breakdown time $t_{*}$ with respect to the number $n_{\rm max}$ of explicitly know Lanczos coefficients. Circles, squares and diamonds correspond to pure initial states polarized in $x$, $y$ and $z$ direction, respectively. The dashed horizontal line marks the time $t_{c}$ of the dynamical quantum phase transition. Clearly, $t_{*}$ exhibits a tendency to saturate around $t_{c}$.
Quench in one dimension.  We focus on a paradigmatic nonintegrable one-dimensional spin-$1/2$ chain – the mixed-field Ising model. Its Hamiltonian reads
$H=\sum\nolimits_{j}\sigma^{x}_{j}\sigma^{x}_{j+1}+h_{z}\sum\nolimits_{j}\sigma^{z}_{j}+h_{x}\sum\nolimits_{j}\sigma^{x}_{j},$ (7)
where $\sigma^{x,y,z}_{j}$ are Pauli matrices at the $j$’th site and $h_{x}$, $h_{z}$ are two parameters of the Hamiltonian. This model is integrable when $h_{x}=0$ or $h_{z}=0$, and nonintegrable otherwise.
The observable we consider here and throughout the paper is the magnetization in the $z$-direction,
$A=\sum\nolimits_{j}\sigma^{z}_{j}.$ (8)
Throughout the paper we consider a family of product translation-invariant initial states with polarization ${\bf p}$,
$\rho_{0}=\bigotimes_{j}\frac{1}{2}(\mathbb{1}+{\bf p}\bm{\sigma}_{j}),$ (9)
where ${\bf p}\bm{\sigma}_{j}=p_{x}\sigma^{x}_{j}+p_{y}\sigma^{y}_{j}+p_{z}\sigma^{z}_{j}$. In general, $|{\bf p}|\leq 1$; in particular, $|{\bf p}|=1$ for pure states with $\rho_{0}=|\Psi_{0}\rangle\langle\Psi_{0}|$.
We are able to compute $n_{\rm max}=45$ nested commutators and Lanczos coefficients. The time evolution of $\langle A_{t}\rangle$ computed by means of the recursion method is presented in Fig. 1. It is benchmarked by the results of exact diagonalization (performed with QuSpin Weinberg and Bukov (2017)) and time-dependent variational principle empowered by matrix product states (performed with ITensor Fishman et al. (2022a, b)). In one dimension the latter methods are essentially exact (and, of course, produce the same result) up to large times, therefore we regard their outcome as actual $\langle A_{t}\rangle$.
One can see from Fig. 1 that the recursion method excellently reproduces the actual $\langle A_{t}\rangle$ at initial time, but then abruptly breaks down. We observe that this breakdown is associated with a dynamical quantum phase transition (DQPT) Pollmann et al. (2010); Heyl et al. (2013); Karrasch and Schuricht (2013); Heyl (2018) – a phenomenon by which the Loschmidt rate function $\lambda(t)=N^{-1}\,\log\big{|}\langle\Psi_{0}|e^{-i\,t\,H}\Psi_{0}\rangle\big{|}^{2}$ features singularities in the thermodynamic limit. Namely, the breakdown time $t_{*}$ is very close tho the time $t_{c}$ of one of the singularities (specifically, the second singularity), see Fig. 1.
The breakdown time $t_{*}$ grows with the number $n_{\rm max}$ of available explicit Lanczos coefficient. However, this growth is saturated at $t_{c}$, as shown in Fig. 2. This strengthens our belief that the observed association between the breakdown of the recursion method and the DQPT is not accidental.
Remarkably, while the recursion method breaks down at a finite time $t_{*}$, at large times it seems to typically converge to the actual equilibrium value of the observable, as illustrated in Fig. 3 (a). Note that the one-dimensional Ising model is known for its atypical thermalization properties Bañuls et al. (2011); Lin and Motrunich (2017); Peng and Cui (2022); Pathak (2024), including long-lived Lin and Motrunich (2017) or even eternal Bañuls et al. (2011) oscillations of polarization around its time-averaged value seen in Fig. 3 (a).
Figure 3: Quench dynamics on (a) one-dimensional nearest-neighbor, (b) square, and (c) cubic lattices. Shown is the expectation value of the total magnetization in the $z$-direction. The initial state is the pure product state (9) polarized along the same direction. Blue solid line is the result of the recursion method. The chosen time span is sufficient for the latter to equilibrate. (a) One-dimensional Ising model (7) with $h_{x}=h_{z}=1$ (the same as in Fig. 1 but for a larger time span). Orange is the exact numerical result. The horizontal dashdotted line is the numerically time-averaged post-quench value of the observable. (b) and (c): The transverse-filed Ising model (10) with $h_{z}=1$ on the square and cubic lattices, respectively. Points are the results of the sparse Pauli dynamics method Begušić and Chan (2024).
Quench in higher dimensions.  In two and three dimensions we consider the transverse-field Ising model on the square and cubic lattices, respectively. The Hamiltonian of the transverse-field Ising model can be written as
$H=\sum_{\langle{\mathbf{i}}{\mathbf{j}}\rangle}\sigma^{x}_{{\mathbf{i}}}\sigma^{x}_{{\mathbf{j}}}+h_{z}\sum_{{\mathbf{j}}}\sigma^{z}_{{\mathbf{j}}},$ (10)
where ${\mathbf{i}}$ and ${\mathbf{j}}$ enumerate lattice sites and the first sum runs over pairs of neighbouring sites.
We compute $22$ and $12$ nested commutators, respectively. Some results are presented in Fig. 3 (b),(c). They are compared to the outcomes of the sparse Pauli dynamics method Begušić and Chan (2024); Begušić .
One can see that the agreement is good initially but abruptly worsens at some threshold time. We are not able to decisively resolve the origin of this discrepancy at the moment. Of course, the DQPT is again the main suspect. DQPT is known to occur in the 2D Ising model, although with a softer logarithmic singularity Heyl (2018). However, we are not able to compute the critical time for this DQPT for two- and three-dimensional models.
Conclusions.  In this work, we have explored the applicability of the recursion method to computing the time evolution of observables relaxing from out-of-equilibrium initial states. We have applied the method to quantum spin systems in one, two, and three dimensions. In one dimension, benchmarking is straightforward due to the availability of essentially exact numerical simulations. By comparing the recursion method to these simulations, we observed that it yields accurate results up to a certain threshold time, beyond which it abruptly breaks down. Notably, we found that this breakdown time closely coincides with the critical time of a dynamical quantum phase transition (DQPT), suggesting that DQPT fundamentally limits the applicability of the recursion method, at least in one dimension.
The underlying mechanism linking DQPT to the breakdown of the recursion method remains entirely unknown and represents a pressing open question. The potential subtlety of this mechanism is highlighted by our observation that the breakdown time coincides not with the first, but with the second (in time) dynamical critical point, implying that not every DQPT necessarily disrupts the recursion method.
In contrast to the one-dimensional case, benchmark data in two and three dimensions is scarce. We have compared our results with those obtained from the sparse Pauli dynamics method Begušić and Chan (2024); Begušić and found that the two approaches diverge abruptly after a certain threshold time, similar to what is observed in one dimension. However, due to the lack of available data on DQPTs in the higher-dimensional models we considered, we cannot unambiguously attribute this divergence to DQPTs. Determining whether DQPTs fundamentally constrain the recursion method’s convergence in higher dimensions remains an open question.
Finally, in one dimension, we have observed that at long times, the recursion method’s outcome coincides with the time-averaged value of the observable, which, in thermalizing systems, is equal to the equilibrium value. If this observation proves to be correct and generic, it would imply an unusual scenario: the recursion method remains valid at both short and long times but fails at intermediate times. Investigating this phenomenon further presents another promising direction for future research.
Note added.  While this manuscript has been in preparation, a closely related study has appeared Loizeau et al. (2025b). The latter paper addresses quench dynamics in one-dimensional spin systems by means of the recursion method. It does not report neither the finite convergence domain of the method nor an association with the DQPT.
Acknowledgements.
Acknowledgments.  We sincerely thank Alexey Rubtsov for pointing to the dynamical phase transition as a potential cause of the recursion method breakdown. We are grateful to Tomislav Begušić and Garnet Kin-Lic Chan for discussing their paper Begušić and Chan (2024), and particularly for providing unpublished data Begušić on the 2D Ising model. Numerous discussions with Alexander Teretenkov and Nikolay Il’in are acknowledged. This work was supported by the Russian Science Foundation under grant № 24-22-00331, https://rscf.ru/en/project/24-22-00331/