### v3

# Quench dynamics via recursion method

Abstract
The recursion method provides a powerful framework for studying quantum many-body dynamics in the Lanczos basis recursively constructed within the Krylov space of operators. Recently, it has been demonstrated that the recursion method, when supplemented by the universal operator growth hypothesis, can effectively compute autocorrelation functions and transport coefficients at infinite temperature. We extend the scope of the recursion method to far-from-equilibrium quench dynamics. We apply the method to spin systems in one, two, and three spatial dimensions. In one dimension its usefulness is limited: although it remains accurate at moderate times, it eventually experiences an abrupt breakdown and subsequently yields results that deviate strongly from the true dynamics. In contrast, in two and three dimensions the method proves far more effective, providing a reasonably accurate description of the evolution across all time scales – from the initial transient regime to thermalization.
Introduction.  Addressing quantum dynamics by means of the recursion method has a long history Mori (1965); Dupuis (1967); Haydock (1980); Mattis (1981); Joslin and Gray (1986); Roldan et al. (1986); Viswanath and Müller (2008); Lindner and Auerbach (2010); Nandy et al. (2024). The (Heisenberg version of the) method essentially amounts to constructing an orthonormal Lanczos basis within the Krylov space of an evolving quantum observable and harnessing the simplifications this basis brings about. A new impact has been given to the method by the universal operator growth hypothesis Parker et al. (2019) – a conjecture about a generic form of the evolution generator in the Lanczos basis.
Up to now the practical application of the recursion method to quantum many-body systems has been limited almost exclusively to computing infinite-temperature autocorrelation functions and transport coefficients Parker et al. (2019); Khait et al. (2016); de Souza et al. (2020); Yates et al. (2020a, b); Yuan et al. (2021); Wang et al. (2024); Bartsch et al. (2024); Uskov and Lychkovskiy (2024); Teretenkov et al. (2024); Bhattacharyya et al. (2024); Füllgraf et al. (2025). There the method has demonstrated high efficiency. This efficiency is particularly notable in application to two-dimensional systems Uskov and Lychkovskiy (2024); Teretenkov et al. (2024); Bhattacharyya et al. (2024), where traditional numerical methods often struggle. Another notable advantage of the recursion method is that it allows a straightforward symbolic implementation, when a single computation covers all parameters of the Hamiltonian Uskov and Lychkovskiy (2024); Teretenkov et al. (2024).
Formally, the output of the recursion method is the Heisenberg operator of an observable. This suggests that the scope of the method can be much broader than infinite-temperature correlation functions and should cover, in particular, the out-of-equilibrium dynamics after quantum quenches. Addressing the later in systems of dimension higher than one is known to be highly challenging, with limited body of results obtained to date Schmitt and Heyl (2020); Li et al. (2020); Dziarmaga (2022); Bernier and Agarwal (2025); Sinibaldi et al. (2024); Gravina et al. (2025); Pavešić et al. (2025); Chen et al. (2025); Begušić and Chan (2025); Loizeau et al. (2025a, b); Rudolph et al. (2025). Given the success of the recursion method in computing two-dimensional autocorrelation function, it seems tempting to apply it to quench dynamics.
Here we undertake exactly this attempt. The paper is organized as follows. We start from briefly introducing the recursion method and specifying it for addressing quench dynamics. Then we apply it to a paradigmatic one-dimensional mixed-field Ising model. The results are benchmarked by essentially exact numerical simulations. Then results for the transverse-field Ising model on the square and cubic lattices are presented. We conclude by discussing implications of our results and open questions raised.
Recursion method for quench dynamics.  Consider an observable given by a self-adjoint Shrödinger operator $A$. The same observable in the Heisenberg representation reads $A_{t}=e^{itH}A\,e^{-itH}$, where $H$ is the Hamiltonian of the system. The Heisenberg operator $A_{t}$ evolves within the Krylov space spanned by operators $A,{\mathcal{L}}A,{\mathcal{L}}^{2}A,\dots$, where ${\mathcal{L}}\equiv[H,\bullet]$ is the superoperator referred to as Liouvillian. $A_{t}$ satisfies the Heisenberg equation of motion $\partial_{t}A_{t}=i\,{\mathcal{L}}A_{t}$. Within the recursion method Viswanath and Müller (2008), this equation is addressed in the orthonormal Lanczos basis $\{O^{0},O^{1},O^{2},\dots\}$ constructed as follows: $O^{0}=A/\|A\|$, $A^{1}={\mathcal{L}}O^{0}$, $b_{1}=\|A^{1}\|$, $O^{1}=b_{1}^{-1}\,A^{1}$ and $A^{n}={\cal L}\,O^{n-1}-\,b_{n-1}O^{n-2}$, $b_{n}=\|A^{n}\|$, $O^{n}=b_{n}^{-1}A^{n}$ for $n\geq 2$. Here a scalar product in the space of operators is introduced according to $\big(A|B\big)\equiv{\cal N}\,\operatorname{\mbox{tr}}\big(A^{\dagger}B\big)$, where $\cal N$} is a normalization constant. For systems of $N$ spins $1/2$ and translation-invariant observables considered here it is convenient to choose ${\cal N}=N^{-1}\,2^{-N}$. The scalar product entails the norm $\|A\|=\sqrt{(A|A)}$. The coefficients $b_{n}$ are referred to as Lanczos coefficients. In the Lanczos basis the Liouvillian has a tridiagonal form,
${\mathcal{L}}=\begin{pmatrix}0&b_{1}&0&0&\dots\\[3.00003pt] b_{1}&0&b_{2}&0&\dots\\[3.00003pt] 0&b_{2}&0&b_{3}&\dots\\[3.00003pt] 0&0&b_{3}&0&\dots\\[3.00003pt] \vdots&\vdots&\vdots&\vdots&\ddots\end{pmatrix}$. (1)
The solution of the Heisenberg equation for the observable $A_{t}$ can be expressed as
$A_{t}=\|A\|\,\sum_{n=0}^{\infty}\phi^{n}_{t}\,O^{n},$ (2)
where $\phi^{n}_{t}\equiv(O^{n}|A_{t})/\|A\|$ are the solution of the system of coupled equations, which can be written in the matrix form as
$\partial_{t}\phi_{t}=$ $i\,{\mathcal{L}}\phi_{t}.$ (3)
Here $\phi_{t}=\|\phi^{n}_{t}\|_{n=0,1,2,\dots}$ is a column vector and $\mathcal{L}$} is the matrix (1).
In practice, one can explicitly compute only a finite number $n_{\rm max}$ of Lanczos coefficients. The rest of the coefficients should be extrapolated. The universal operator growth hypothesis Parker et al. (2019) (see also precursor work Liu and Müller (1990); Florencio et al. (1992); Zobov and Lundin (2006); Elsayed et al. (2014); Bouch (2015)) states that in a generic quantum many-body system Lanczos coefficients $b_{n}$ scale asymptotically linearly with $n$ at large $n$ (with a logarithmic correction for one-dimensional systems). This asymptotic behavior has been confirmed in various many-body models Cao (2021); Noh (2021); Heveling et al. (2022); Uskov and Lychkovskiy (2024); De et al. (2024); Ermakov (2025). As a rule, the leading asymptotics can be easily extracted from the first dozen of Lanczos coefficients. However, extrapolating Lanczos coefficients by a linear function is a too rough approximation: the qualitative behavior of the solutions of the Heisenberg equation strongly depends on the subleading asymptotic terms Viswanath et al. (1994); Parker et al. (2019); Yates et al. (2020a, b); Dymarsky and Smolkin (2021); Bhattacharjee et al. (2022); Avdoshkin et al. (2024); Camargo et al. (2023); Uskov and Lychkovskiy (2024); Dodelson (2025); Gamayun et al. (2025). Guided by the results of ref. Gamayun et al. (2025), we employ the following extrapolation formula:
$b_{n}$ $\simeq\frac{\alpha\,n}{\log n}+\gamma+\frac{(-1)^{n}\tilde{\alpha}}{\log n+c}$ ${\rm for}~~1{\rm D},$ (4)
$b_{n}$ $\simeq\alpha\,n+\gamma+\frac{(-1)^{n}\tilde{\alpha}}{\log n+c}$ ${\rm for}~~2,3{\rm D}.$ (5)
The last alternating term is required whenever $\varphi^{0}_{t}$ relaxes to a nonzero value at $t\rightarrow\infty$ Gamayun et al. (2025), which is generally the case. Parameters $\alpha,\,\gamma$ and $c$ are found by fitting $n_{\rm max}$ explicitly computed Lanczos coefficients.
The knowledge of the Heisenberg operator $A_{t}$ of the observable allows one to address the quench dynamics, i.e. compute the time-dependent value $\langle A_{t}\rangle\equiv\operatorname{\mbox{tr}}\rho_{0}A_{t}$ of the observable that evolves after the system is initialized in an out-of-equilibrium state $\rho_{0}$. Within our approach, we compute $\langle A_{t}\rangle$ according to
$\langle A_{t}\rangle\simeq\sum_{n=0}^{n_{\rm max}}\phi^{n}_{t}\,\langle O^{n}\rangle,$ (6)
where $\langle O^{n}\rangle\equiv\operatorname{\mbox{tr}}\rho_{0}O^{n}$. We compute $O^{n}$ and $b_{n}$ for $n$ up to $n_{\rm max}$ symbolically directly in the thermodynamic limit, as described in Uskov and Lychkovskiy (2024). $\langle O^{n}\rangle$ is also computed symbolically. Final numerical values of $b_{n}$ are obtained with the use of rational arithmetic, which avoids the numerical instability of the Lanczos algorithm Eckseler et al. (2025). As for functions $\phi^{n}_{t}$, they are obtained by numerically solving coupled linear differential equations (3) with the tridiagonal matrix $\mathcal{L}$} truncated at some large $k$ (we use $k=1000$ in the examples below). The entries of this matrix for $n>n_{\rm max}$ are obtained from extrapolation of the $n_{\rm max}$ known Lanczos coefficients according to extrapolation formulae (4), (5).
It should be emphasized that the sequence $\langle O^{n}\rangle$ generally does not have any apparent structure and is not expected to have any particular asymptotic behavior, in contrast to the sequence of Lanczos coefficients $b_{n}$. We further discuss this fact and illustrate it by examples in the Supplement sup .
Note that $\phi^{0}_{t}=({\cal N}/\|A\|)\,\operatorname{\mbox{tr}}A\,A_{t}$ is the infinite-temperature autocorrelation function of the observable $A$, up to a constant factor. Recently the recursion method supplemented by the UOGH has proven effective in computing such autocorrelation functions Parker et al. (2019); Wang et al. (2024); Uskov and Lychkovskiy (2024); Teretenkov et al. (2024). The out-of-equilibrium post-quench expectation value $\langle A_{t}\rangle$ is, however, a more complex object. Below we explore it in specific models.
Figure 1: Time evolution of the total polarization along $z$ direction in the 1D Ising model (7) with $h_{x}=h_{z}=1$. The initial state is a product translation-invariant state (9) with the polarization $(p_{x},p_{y},p_{z})$ indicated on each plot. Orange line – essentially exact results of the numerical time evolution (see text for details). Blue line - results of the recursion method. The recursion method reproduces initial evolution but abruptly departs from the exact numerical result at the breakdown time $t_{*}$ (vertical dashed line). This time appears to be independent of the initial state.
Quench in one dimension.  We focus on a paradigmatic nonintegrable one-dimensional spin-$1/2$ chain – the mixed-field Ising model. Its Hamiltonian reads
$H=\sum\nolimits_{j}\sigma^{x}_{j}\sigma^{x}_{j+1}+h_{z}\sum\nolimits_{j}\sigma^{z}_{j}+h_{x}\sum\nolimits_{j}\sigma^{x}_{j},$ (7)
where $\sigma^{x,y,z}_{j}$ are Pauli matrices at the $j$’th site and $h_{x}$, $h_{z}$ are two parameters of the Hamiltonian. This model is integrable when $h_{x}=0$ or $h_{z}=0$, and nonintegrable otherwise.
The observable we consider here and throughout the paper is the magnetization in the $z$-direction,
$A=\sum\nolimits_{j}\sigma^{z}_{j}.$ (8)
Throughout the paper we consider a family of product translation-invariant initial states with polarization $\bf p$},
$\rho_{0}=\bigotimes_{j}\frac{1}{2}(\mathbb{1}+{\bf p}\bm{\sigma}_{j}),$ (9)
where ${\bf p}\bm{\sigma}_{j}=p_{x}\sigma^{x}_{j}+p_{y}\sigma^{y}_{j}+p_{z}\sigma^{z}_{j}$. In general, $|{\bf p}|\leq 1$; in particular, $|{\bf p}|=1$ for pure states with $\rho_{0}=|\Psi_{0}\rangle\langle\Psi_{0}|$.
We are able to compute $n_{\rm max}=48$ nested commutators and Lanczos coefficients for the Hamiltonian (7) and the observable (8). The time evolution of $\langle A_{t}\rangle$ computed by means of the recursion method, i.e. from eq. (6), is presented in Fig. 1. Values of $b_{n}$, as well as individual ingredients entering eq. (6), i.e. $\phi^{n}_{t}$ and $\langle O^{n}\rangle$, are presented in the Supplement sup .
Figure 2: Quench dynamics of nonintegrable Ising models on (a) one-dimensional, (b) square, and (c) cubic lattices. Shown is the expectation value of the total magnetization in the $z$-direction. The initial state is the pure product state (9) polarized along the same $z$ direction. Blue solid line is the result of the recursion method. The chosen time span is sufficient for the latter to equilibrate. (a) One-dimensional Ising model (7) with $h_{x}=h_{z}=1$ (the same as in Fig. 1 but for a larger time span). Orange is the exact numerical result. The horizontal dashdotted line is the numerically time-averaged post-quench value of the observable. (b) and (c): The transverse-filed Ising model (10) with $h_{z}=1$ on the square and cubic lattices, respectively. Points are the results of the sparse Pauli dynamics method Begušić and Chan (2025). Insets show explicitly computed Lanczos coefficients (points) and their extrapolation according to eqs. (4), (5) (black lines).
The results of the recursion method are benchmarked by the results of exact diagonalization performed with QuSpin Weinberg and Bukov (2017) and time-dependent variational principle empowered by matrix product states performed with ITensor Fishman et al. (2022a, b) (see the Supplement sup for details). In one dimension these numerical methods are essentially exact up to sufficiently large times, therefore we regard their outcome as actual $\langle A_{t}\rangle$.
One can see from Fig. 1 that the recursion method excellently reproduces the actual $\langle A_{t}\rangle$ initially, but then abruptly breaks down at the breakdown time $t=t_{*}$. After this breakdown, the recursion method result strongly deviates from the actual value of the observable. We note that the breakdown time appears to be independent of the initial state.
We attribute this breakdown of the recursion method to the slow thermalization typical of one-dimensional systems. In particular, the mixed-field Ising model is well known for its anomalous thermalization behavior Bañuls et al. (2011); Lin and Motrunich (2017); Kormos et al. (2017); Peng and Cui (2022); Pathak (2024).
In the previous version of this manuscript Shirokov et al. (2025), we linked the breakdown of the recursion method to a dynamical quantum phase transition (DQPT), motivated by an apparent coincidence between the breakdown time and the DQPT time. A more careful analysis shows that this coincidence is not universal and does not hold across most of the parameter space. We therefore retract our earlier claim that a DQPT is the cause of, or a universal precursor to, the breakdown of the recursion method in one dimension. Additional data and discussion on this point are provided in the Supplement.
Remarkably, at large times the recursion method result typically converges to the actual equilibrium value of the observable, as illustrated in Fig. 2 (a). This convergence may indicate that the terms discarded by the truncation in Eq. (2), while significantly affecting the transient dynamics, do not contribute much in the limit of $t\rightarrow\infty$.
Quench in higher dimensions.  In two and three dimensions we consider the transverse-field Ising model on the square and cubic lattices, respectively. The Hamiltonian of the transverse-field Ising model can be written as
$H=\sum_{\langle{\mathbf{i}}{\mathbf{j}}\rangle}\sigma^{x}_{{\mathbf{i}}}\sigma^{x}_{{\mathbf{j}}}+h_{z}\sum_{{\mathbf{j}}}\sigma^{z}_{{\mathbf{j}}},$ (10)
where $\mathbf{i}$} and $\mathbf{j}$} enumerate lattice sites and the first sum runs over pairs of neighbouring sites. The observable is the same as in the previous case – the total magnetization in the $z$-direction, $A=\sum_{{\mathbf{j}}}\sigma^{z}_{{\mathbf{j}}}$.
We compute $23$ and $17$ nested commutators, respectively. Time evolution of the magnetization is presented in Fig. 2 (b),(c). The results of the recursion method are compared to the outcomes of the sparse Pauli dynamics (SPD) method Begušić and Chan (2025); Beg . Up to the maximal time accessible to the SPD, the agreement with our results is excellent at early times and remains reasonably good thereafter. In contrast to the SPD, however, the recursion method yields predictions all the way to equilibration.
Conclusions and discussion.  In this work, we have explored the applicability of the recursion method to computing the time evolution of observables relaxing from out-of-equilibrium initial states. We have applied the method to quantum spin systems in one, two, and three dimensions.
In one dimension, benchmarking is straightforward due to the availability of essentially exact numerical simulations. By comparing the recursion method to these simulations, we have observed that it yields accurate results up to a certain threshold time, beyond which it abruptly breaks down. Despite this breakdown, the recursion method result eventually equilibrates close to the actual equilibrium value.
The breakdown indicates the slow convergence of the expansion (6) of a Heisenberg observable in the Lanczos basis. This slow convergence may limit the scope of concepts and techniques Parker et al. (2019); Chu et al. (2024); Loizeau et al. (2025c); Alishahiha and Vasli (2025); Rabinovici et al. (2025) built on this expansion in one dimension.
In contrast to the one-dimensional case, benchmark data in two – and especially three – dimensions remain scarce. We have compared our results with those obtained using the SPD method Begušić and Chan (2025); Beg and find agreement ranging from excellent to reasonable up to the maximal times accessible to SPD. Importantly, the recursion method is essentially unrestricted in time and continues to yield reasonable results all the way to equilibration.
Note added.  While this manuscript has been in preparation, a closely related study has appeared Loizeau et al. (2025c). The latter paper addresses quench dynamics in one-dimensional spin systems by means of the recursion method.
Acknowledgements.
Acknowledgments.  We are grateful to Tomislav Begušić and Garnet Kin-Lic Chan for discussing their paper Begušić and Chan (2025) and our manuscript, and particularly for providing unpublished data Beg on the 2D Ising model. We thank Berislav Buča and Nicolas Loizeau for useful remarks and discussions. Finally, numerous discussions with Alexander Teretenkov and Nikolay Il’in are gratefully acknowledged. This work was supported by the Russian Science Foundation under grant № 24-22-00331, https://rscf.ru/en/project/24-22-00331/
Appendix A Supplementary material
Appendix B S1. Details on $\phi^{n}_{t}$ and $O^{n}_{t}$
For a Hermitian operator $A$, the operator ${\mathcal{L}}^{n}A$ is Hermitian for even $n$ and anti-Hermitian for odd $n$. Therefore $O^{n}$ is Hermitian (anti-hermitian) for even (odd) $n$, and $\phi^{n}_{t}$ are real (imaginary) for even (odd) $n$. This $(-i)^{n}\phi^{n}_{t}$ and $i^{n}\langle O^{n}\rangle$ are always real. We plot these quantities in Fig. S1 and Fig. S2, respectively.
Figure S1: Functions $(-i)^{n}\phi^{n}_{t}$ for $n=0,1,\dots,n_{\rm max}$. Shown are plots for (a) one-dimensional, (b) two-dimensional and (c) three-dimensional Ising models considered in the main text.
As noted in the main text, $i^{n}\langle O^{n}\rangle$ do not exhibit any simple structure, as can be seen in Fig. S2. Importantly, available data show that convergence of $\big|\langle O^{n}\rangle\big|$ is either absent or, at least, rather slow. In fact, we have no reasons to expect fast convergence. Indeed, a straightforward sum rule for these coefficients,
$\sum_{n=0}^{\infty}\big|\langle O^{n}\rangle\big|^{2}=\sum_{n=0}^{\infty}\left(N\,2^{N}\right)^{2}\,(\rho_{0}|O^{n})(O^{n}|\rho_{0})=\left(N\,2^{N}\right)^{2}\,(\rho_{0}|\rho_{0})=N\,2^{N}\operatorname{\mbox{tr}}\rho_{0}^{2}\geq N$ (S1)
diverges in the thermodynamic limit of $N\rightarrow\infty$. This implies that $\big|\langle O^{n}\rangle\big|$ can not converge to zero faster than $1/\sqrt{n}$. It remains an interesting open question whether this slow (or absent) convergence is responsible for the breakdown of the recursion method.
Figure S2: Values of $i^{n}\langle O^{n}\rangle$ for the one-dimensional (left column) and two-dimensional (right column) Ising models considered in the main text. Polarizations of the initial state are indicated in each plot.
Appendix C S2. Numerics in the one-dimensional case
For a one-dimensional model (7), we benchmark the time evolution of the observable by numerical methods applied to finite systems. Specifically, we perform exact diagonalization for up to $N=14$ spins with QuSpin Weinberg and Bukov (2017) and matrix product state-empowered time-dependent variational (TDVP) computation for up to $N=40$ spins. All numerical results coincide without visible discrepancies up to at least $t=3.5$, thus covering the time span of Fig. 1. For larger times shown in Fig. 2(a) we use the TDVP result for $N=40$.
Appendix D S3. Moments and Lanczos coefficients
Below we provide values of moments $\mu_{2n}=(A|{\mathcal{L}}^{2n}A)/(A|A)$ and Lanczos coefficients $b_{n}$ for models considered in the main text. An algorithm to obtain the latter from the former can be found in Dupuis (1967); Viswanath and Müller (2008). Note that, for technical reasons, in the 3D case we compute moments and Lanczos coefficients not for a translation-invariant observable (8), as for 1D and 2D cases, but for its local counterpart $\sigma^{z}_{{\mathbf{j}}}$. While moments and Lanczos coefficients of these two observables are different, their time dependent expectation values are equal up to an overall factor $N$ as long as the initial state is translation-invariant, thanks to the translation invariance of the Hamiltonian.
Note that in the previous version of the manuscript Shirokov et al. (2025), for the 3D Ising model, we provided 12 moments and Lanczos coefficients for the non-translation-invariant observable $\sigma^{z}_{j}$. Here we provide 17 moments and Lanczos coefficients for the respective translation-invariant observable. While the values are different in the two cases, the output of the recursion method is consistent whenever the initial state is translation-invariant.
$n$ $\mu_{2n}$ $b_{n}$
1 12 3.4641
2 480 5.2915
3 25984 4.4934
4 1694208 6.0281
5 127258624 5.6190
6 10783342592 6.9438
7 1019673509888 6.6972
8 107060323680256 8.2286
9 12469034121428992 7.9953
10 1613762515315982336 9.2837
11 232736621970029805568 8.9148
12 37468646645944075419648 10.2773
13 6726394693098760675786752 9.8677
14 1340433610285320688734568448 11.1094
15 294391435294882867890188451840 10.7766
16 70659698635880317667608199430144 12.1889
17 18383424921170050832257258778263552 11.7912
18 5147501922089241875609891362536685568 12.9532
19 1542402400283276113669614948668265201664 12.7320
20 492432525422001580271341582388177351475200 13.9288
21 166985805198508106978965686562725967100379136 13.5778
22 60011852488095956067395190665735924993040056320 14.6867
23 22821321744324019152247154618657381361730948431872 14.4673
24 9172207374041781749043802673339425314176975748726784 15.5798
25 3892052853514868229060585022861205953766446860739805184 15.3892
26 1741736488569588529172771998235278437950082051065922977792 16.4064
27 821019133269536743838893539994338673397100678325485105577984 16.1096
28 407086843763398331397385204208407255958886525287040747053252608 17.3132
29 211990829889808789195437699387039831255135134538810758197371994112 17.0171
30 115754651941408887232253140259122456149371760683740727687331373383680 18.0978
31 66166596237703872256126915723522220984465097504279554678985792092635136 17.7211
32 39530072936257449901374525090449349559603851447100658065034700516313006080 18.8846
33 24646757783977889749525872721004379593240131044109640793744075941140506869760 18.6810
34 16015858109930051337482877843244494350822380543617937802711923518036880757096448 19.6477
35 10833753072406488462262074132859724923839003624262790509360495364480719760857759744 19.4833
36 7620581541304124813656489403945273131158245029661872172588240591095613039962778238976 20.2878
37 5568944585532111508410499104658395233823379770020424361966461875062472358794201533513728 20.3901
38 4224439735905136953825989317559917466695928508761395026089245710815314351095999194227277824 21.0909
39 3323843247775625289903213303600983380005695346567953993918161657219885198513483978643486015488 21.2234
40 2710613205308213119249609362155540213052003237873428947613674131447949025132969096999139593420800 21.7286
41 2289489892850970408834540810440144834655617422363836454740677715045134451758621541567258816170426368 21.9363
42 2001453495607268189063620952290269505902583051218882018390254310106562921877530480766976590064753049600 22.6224
43 1809588827470560013981821635403092794798688775869840545515811131487657718540062168703067251339768165826560 22.7147
Table S1: Moments and Lanczos coefficients for the 1D Ising model (7) at $h_{x}=h_{z}=1$ and observable (8). Moments for $n=44,\dots,48$ are too large to fit the page. They can be obtained from the authors upon a request.
$n$ $\mu_{2n}$ $b_{n}$
1 16 4.0000
2 896 6.3246
3 71680 5.7966
4 7307264 7.9714
5 908099584 7.7771
6 134882656256 9.8932
7 23708042264576 9.8014
8 4897950853496832 11.7753
9 1181552437773729792 11.8055
10 330379916698579894272 13.6724
11 106222899499818563928064 13.7546
12 38952574279318656784531456 15.6436
13 16166117065306264371967557632 15.7188
14 7539715274735761761366362292224 17.3406
15 3926354218204565053083857654382592 17.3814
16 2269213981250751881428716107876794368 19.0685
17 1446793491966272434190827088127549505536 19.2419
18 1011459407804179014361425714300323771711488 20.8675
19 770720014993499216420856132351984701670424576 21.0498
20 636516730341979806791563685949386538476858507264 22.6354
21 566952387953410930810990384866974059612775307542528 22.8705
22 542434116478048005253046423594975291775134675528318976 24.4221
23 555699452970039450718350263160732485341995086705577689088 24.6588
Table S2: Moments and Lanczos coefficients for the observable (8) and the Ising model on the square lattice, eq. (10), with $h_{z}=1$.
$n$ $\mu_{2n}$ $b_{n}$
1 24 4.89898
2 1920 7.48331
3 230400 7.55929
4 36618240 9.98284
5 7313326080 10.1916
6 1790621122560 12.7955
7 530848042450944 13.1431
8 188884569821282304 15.4032
9 79892590021627084800 15.8832
10 39721461646806598287360 18.2052
11 22947716425271653039079424 18.7155
12 15244534918363394241158184960 20.9982
13 11543604835131049187075722051584 21.5619
14 9891579029630566720438258058133504 23.8488
15 9532138756054379075322538513856987136 24.4597
16 10273436716844388765436471742460667625472 26.751
17 12321137907981191488465349016456502941057024 27.4636
Table S3: Moments and Lanczos coefficients for the observable $\sigma^{z}_{{\mathbf{j}}}$ and the Ising model on the cubic lattice, eq. (10), with $h_{z}=1$.
Appendix E S4. Recursion method breakdown and dynamical quantum phase transition in one dimension
Figure S3: Time evolution of the total polarization along $z$ direction in the 1D Ising model (7) for various values of $h_{z}$ and $h_{x}$. The initial state is a product translation-invariant state (9) polarized in the $z$-direction. Orange line – essentially exact results of the numerical time evolution (see text for details). Blue line – results of the recursion method. Pink line – the rate function $\lambda(t)$ whose cusps mark the dynamical quantum phase transitions.
In the previous version of this manuscript Shirokov et al. (2025), we have attributed the breakdown of the recursion method to a dynamical quantum phase transition (DQPT) Pollmann et al. (2010); Heyl et al. (2013); Karrasch and Schuricht (2013); Heyl (2018) – a phenomenon by which the Loschmidt rate function $\lambda(t)=-N^{-1}\,\log\big|\langle\Psi_{0}|e^{-i\,t\,H}|\Psi_{0}\rangle\big|^{2}$ features singularities in the thermodynamic limit. [^1] There is an ambiguity in the meaning of the term “dynamical (quantum) phase transition.” Historically, it was introduced in a quite different context, referring to a transition between distinct phases of matter that unfolds over time due to variations in the Hamiltonian Schützhold et al. (2006); Eckstein et al. (2009); Uhlmann et al. (2010); Garrahan and Lesanovsky (2010); Diehl et al. (2010); Sciolla and Biroli (2010) – a definition more closely aligned with the conventional understanding of phase transitions. In contrast, the definition we adopt was introduced later Heyl et al. (2013) and is based on a formal analogy between the Loschmidt rate and the free energy.
Here, we have examined this conjecture more thoroughly across a range of model parameters. The results, shown in Fig. S3, indicate that, in general, $t_{*}$ and $t_{c}$ do not coincide. The apparent agreement occurs, within uncertainties, only for $h_{x}\simeq h_{z}$. We therefore retract our earlier strong claim regarding a relation between the recursion method breakdown and the DQPT, leaving further investigation of this connection to future work. For the sake of this future work, we point to a paper Takahashi (2025) on DQPT and Lanczos method in the Schrödinger representation.